

********************************************************************
UARS WEEKLY STATUS REPORT - 09 JUL 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM gathered its normal solar, offset, and occultation data this
past week without apparent problems. Evaluation of the recently
reprocessed V20 products before their release to the DAAC continues.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
HALOE was turned off at the end-of-day July 6, 1999 in preparation for the
YAW. HALOE is scheduled to turn on July 14, 1999.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Owing to power limitations, it was not possible to activate WINDII
for the June 30 - July 9 operating period as planned.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF notified all Instrument PI Software Reps regarding data gap for
UARS DAY 2849 @13:21:35.06 - 14:01:59.60.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
18 days CLA (7.0) PROD (L1-L3) (475R-89C-11F=375L)
57 days WIN (5.11) PROD (L1-L3) (1540R-894C-7F=639L)
The CDHF will load May 1999 current day DAAC data in July 1999.
CDHF pending software upgrades: DEC AXP Fortran 6.5, DEC AXP OPEN
VMS 6.2, OpenIngres.
***********
FOT
***********
REPORTING PERIOD
This report covers 26 June 1999 (Orbit 42563 , GMT Day 177) through 02 July
GMT Day 183).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
MLS 180/2306 42622 Off
ACRIM 181/0007 42623 On
SUSIM 181/0007 42623 On
PEM 181/0646 42627 AXIS On
WINDII 181/1620 42633 On
WINDII 181/1807 42634 Off
WINDII 181/2253 42637 On
WINDII 182/0015 42638 Off
The beta angle ranged between the angles of 33.5 degrees to 26.0 degrees
for this report period. The beta angle is now decreasing toward a minimum
angle of 2.0 degrees on 10 July 1999 (DOY 191).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.7 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.7 to 1.0 degrees C. Additional battery performance data
for each battery on 25 June 1999 and 02 July 1999 are provided in Appendix
D.
The clock error ranged from 9.7 to -11.9 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% WINDII 06
% FOT 04
% SUSIM 02
% MLS 02
% ACRIM 01
% PEM 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 02
Three (3) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 99/130 - 2287.496913
Transponder B: 99/151 - 2287.499087
GROUND SYSTEM OPERATION
Two (2) operational support problem occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 99/119 show the observatory data loss to be
24 hours, 53 minutes, 57.044 seconds (an increase of 40 minutes and 48.704
seconds since the last report period). This is a 0.0365 percent data loss
which equals a 99.9635 percent data capture for the mission.
The increase of 40 minutes and 48.704 seconds occurred over one (1) day:
7 40 minutes and 48.704 seconds on Day 181 due to WSGT hardware
problems.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
FOT personnel are supporting the IMOC (SCS-21) development effort. The
current work is a continuation of the SCS-21 UARS implementation effort
done last year which culminated in a level 1 demonstration.
The FOT continues to investigate the feasibility of a new PMON action to
change battery V/T curves based on loading to mitigate battery risks in sun
pointing safehold.
YEAR 2000 OPERABILITY SUPPORT
The end-to-end testing has been delayed until a new FORMATS software
release can be made to accommodate the special Y2K planning aids.
MEETINGS
The FOT met with the IMOC Systems Engineering Team to explain some of the
UARS requirements.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver (reverse-to-forward) is scheduled for 10
July 1999 (DOY 191).
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 02 July 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
42563 177/0018 99-219 OBC Clk Adj 0.00 to 20.6 msec/day
42585 178/1207 99-220 OBC Clk Adj fr 20.6 to 0.00 msec/day
42613 180/0840 99-221 MLS MLS Scan Stop/Run
42622 180/2306 99-222 MLS Power OFF
42623 181/0007 99-223 ACRIM Power ON
42623 181/0007 99-224 SUSIM Power ON
42627 181/0646 99-225 PEM AXIS Power ON
42633 181/1620 99-226 WINDII Power ON
42634 181/1807 99-227 WINDII Power OFF
42636 181/2053 99-228 OBC Clk Adj 0.00 to 20.6 msec/day
42637 181/2253 99-226 WINDII Power ON
42638 182/0015 99-229 WINDII Power OFF
42643 182/0852 99-230 WINDII PSU-A OFF
42644 182/1035 99-231 WINDII SURVIVAL HTR A ON
42649 182/1811 99-232 SUSIM DETECTOR OFF
42655 183/0418 99-223 OBC Clk Adj fr 20.6 to 0.00 msec/day
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 02 July 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 02 July 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
42563 177/00:27 (25 June 99) from 00.0 to 20.6 msec/day
42585 178/12:09 (27 June 99) from 20.6 to 00.0 msec/day
42636 181/20:56 (30 June 99) from 00.0 to 20.6 msec/day
42655 183/04:24 (02 July 99) from 20.6 to 00.0 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 02 July 1999
25 June 1999 - GMT Day 176 Beta = 33.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -95.2/-123.2 -2.42/-3.68 +0.0/ +0.0 16.5 0.000 0.0
2 +22.4/-39.2 +6.44/+5.47 +29.2/-15.2 24.2 1.023 22.5
3 +22.4/-67.2 +5.47/+4.51 +28.8/-15.6 24.2 1.025 22.2
02 July 1999 - GMT Day 183 Beta = 26.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -100.8/-123.2 -3.05/-4.31 +0.0/ +0.0 16.5 0.000 0.0
2 +11.2/-72.8 +6.11/+4.51 +29.6/-16.0 23.8 1.023 21.5
3 +22.4/-56.0 +5.15/+3.87 +28.8/-15.6 23.7 1.020 21.2
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 02 July 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-031 FOT 42629 181/1033 21028 NEG ACQ. 15:00 RECOV.
99-032 FOT 42629 181/1220 21028 NEG ACQ 00:41:04 LOST
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-028 FOT see list June 99 14170 Generic Late Acq.
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-033 FOT see list July 99 14170 Generic Late Acq.
99-031 FOT 42629 181/1033 21028 NEG ACQ. 15:00 RECOV.
99-032 FOT 42629 181/1220 21028 NEG ACQ 00:41:04 LOST
SEVNTFW99/27